• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Saha

    Log entry time 01:48:13 on June29,2002

    Entry number 87011

    keyword=Owl shift summary - Pb target test

    00:00 Taking data at point Pb_B at 25 uA
    00:18 1900 run finished (1.3mln events), changed beam to 45uA.
    00:26 roc14 crashed
    00:31 1902 started, 45 uA
    00:40 ...ended, 1M events
    1903 started, 45 uA
    0050: 1903 ended, 1M events
    0053: 1904 started, 15 uA
    0058: 1904 ended, CODA crashed, 100k events
    0105: 1905 started, 60 uA
    0110: 1905 crashed, ROC14, 200k events
    dropped current to 15 uA.
    0115: 1906 started, 15 uA
    0120 1906 ended, 77k events, Lead target still there.
    going to 75 uA
    0125: 1907 started, 75 uA
    0133: 1907 ended, 291k events, middle foil seems to be missing
    0134: 1908 started, 15 uA
    It seems only one foil is left.
    0143: 1908 ended, 108k events.

    -- End of lead target test, Only one foil remaining.